Hello, this is my math question relating to number sequences. Is there a specific equation for a number plus(+) every whole number less than it down to 1? [Ex.] 365+364+363+...1=?
1
Expert's answer
2013-01-11T06:15:48-0500
A sequence such as 1, 2, ... 365 or 365, 364, .. 1 which has aconstant difference between terms called arithmetic progression. The first
term is a1, the common difference is d, and the number of terms is n.
an = a1 + (n – 1)d
the sum of an arithmeticprogression is equal to
Sn = (n/2)(a1+an)
or
Sn = n/2 * (2*a1+d(n-1))
S(365)=365/2*(1+365)= 66795